OpenStudy (anonymous):

(256^0.16*256^0.09)^2=

ganeshie8 (ganeshie8):

\(\large \left(256^{0.16}*256^{0.09}\right)^2 \) \(\large \left(256^{0.16+0.09}\right)^2 \)

ganeshie8 (ganeshie8):

add the exponents

ganeshie8 (ganeshie8):

it simplifies nicely

OpenStudy (anonymous):

(256^0.25)^2

ganeshie8 (ganeshie8):

yes, next recall that 0.25 is just one quarter : 0.25 = 1/4

OpenStudy (anonymous):

16

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

i have another question